Sunni’s retail furniture store has bought new sofas from its supplier at a cost of $800.00 each.  Sunni intends to sell each new sofa at a mark-up of 130%, based on cost.   What is the rate of markup based on the selling price?
Answer #1

Mark up Amount = 800*130% = $1,040 Selling price = cost + Mark up = 800+1040 = $1,840 Mark up on selling price = 1040/1840 =